# A theoretical analysis on the resolution of parametric PDEs via neural networks designed with Strassen algorithm

### v1

@@ Abstract:
We construct a Neural Network that approximates the matrix multiplication operator for any activation functions for which there exist a Neural Network which can approximate the scalar multiplication function. In particular, we use the Strassen algorithm for reducing the number of weights and layers needed for such Neural Network. This allows us to define another Neural Network for approximating the inverse matrix operator. Also, by relying on the Galerkin method, we apply those Neural Networks for resolving parametric elliptic PDEs for a whole set of parameters at the same time. Finally, we discuss the improvements with respect to prior results.
@@ 1 Introduction
The Machine Learning explosion has left little research branches untouched. Numerical Analysis for PDEs is no exception. The Physics Informed Neural Network is an example of this: using Neural Networks as approximate solution of a particular PDE [RPK19]. However, this is not the only example, in [KPRS22], for instance, the authors construct two $\operatorname{ReLU}$ Neural Networks capable to approximate the matrix multiplication and matrix inverse operators while giving upper bound on the number of weights and layers. These Neural Networks are then used in the Galerkin method for resolving elliptic PDEs. The cost of using these methods over conventional ones is hard to justify for solving a single particular PDE. However, they are useful for numerically solving parametric PDEs for a set of parameters.
The use of Machine Learning for parametric PDEs tries to avoid the problems of solving them for each parameter. These problems are essentially the study of too many parameters or a high computing complexity for even a single PDE which both leads to an unreasonable time cost. When solving for a set of parameters at the same time, those problems are mitigated.
Parametric elliptic PDEs appear in many contexts such as physics and engineering. The problem of describing a physical phenomenon sometimes translates into solving an elliptic PDE. Many examples of this exist, such as the stationary Navier-Stokes, Poisson, Laplace, Helmholtz, and anisotropic heat diffusion equations, [Pol01, PZ03]. Those equations explains fluid mechanics, electromagnetism and gravitational fields, and steady-state processes of heat/mass transfers respectively. Given its frequency and importance, there is a need to numerically solve these equations effectively.
@@ 1.1 Our Contribution
Motivated by the results of [KPRS22], we search to understand neural networks while improving the bounds on the number of parameters via mathematical arguments. Our contributions are the following ones:
@@
  • Activation functions. In Neural Networks, the activation function plays a major role, as it determines its expressivity. In this paper, we restrict ourselves to activation functions for which there are Neural Networks that approximates the scalar product operation, as mathematically stated in Hypothesis 1. Clearly, if we can not approximate such operation, there is no hope of representing the product of matrices. This set of maps contains $\operatorname{ReLU}$, the only activation function considered in [KPRS22].
@@
  • Bounds. The authors of [KPRS22] bound the number of parameters needed for neural networks used for parametric PDEs. To achieve better bounds, we use two key ideas. The first one is to use the Strassen algorithm for matrix multiplication, which algorithm reduces the number of scalar products needed for matrix multiplication. This allows us to economize the use of the potentially bulky Neural Network that approximates the product of scalars. The second one is the use of a finite sum of powers of matrices to approximate the inverses of matrices appearing in the Galerkin method. Rewriting cleverly that sum, we reduce the number of matrix multiplications in comparison to [KPRS22].
@@
    In paricular, under the hypothesis on the activation function, we prove two main theorems.
@@
    The first one, Theorem 3.3, is a construction of a Neural Network that approximates the matrix multiplication operator. We achieve a smaller size of Neural Network than the one in [KPRS22] where its dependency on the number of rows of the input matrices is no longer cubic, but of the order of $\log_{2}7$.
@@
    The second one, in Theorem 4.6, we get the same result, but for approximating the matrix inverse operator. In Section 5, we see how Theorem 4.6 has an use in numerical solving of PDEs with the Galerkin method. The most common application of the Galerkin method, the finite element method, suffers from the curse of dimensionality which we diminish with the exponent $\log_{2}7$. Finally, in Section 6, we compare both theorems with those in [KPRS22] and show the improvements.
@@
  • Notation. Finally, we simplify notation by allowing matrices as input of Neural Networks. This avoids flattening matrices which overloads the notation.
1.2 Related Work
@@
In this subsection, we discuss the results in the literature related to this paper. The most related work is [KPRS22] since this work is a generalization, and improvement of its results, as we have explained before.
@@
For the rest of the literature review, first in Section 1.2.1, we explore multiplication matrix algorithm which optimizes the amount of scalar products. We follow in Section 1.2.2 by presenting approximation theory for Neural Networks, which justifies that the product of scalars can be approximated by Neural Networks with a great variety of activation functions. Then, in Section 1.2.3, we explore the different interactions involving matrix multiplication, Neural Networks and the Strassen algorithm. Finally, in Section 1.2.4, we see the uses that Neural Networks receive to numerically solve PDEs.
@@ 1.2.1 Strassen-Like Algorithms
In [Str69], the optimality of matrix multiplication is questioned when the author showed that for matrices of size $n\times n$, it is only required $\mathcal{O}(n^{\log_{2}7})$ scalar products.
Clearly, one would like to find the infimum of $\omega$ such that for every $\varepsilon>0$ there is a $(n\times n)$-matrix multiplication algorithm with $\mathcal{O}(n^{\omega+\varepsilon})$ arithmetic operations. For that purpose a number of more optimal algorithms than Strassen’s appeared. For example, we would like to highlight two of them: [CW90] describes an algorithm that was not outperformed for 20 years, and the latest and lowest upper bound at the moment enounced in [ADW+24]. Both use the laser method introduced by Strassen in [Str87] that relates trilinear polynomials to the matrix multiplication tensor. This method, however, has its limitation: [AFLG15] shows that using that method and a variety of its generalizations can not achieve the target $\omega=2$.
Even if those papers present better algorithms than the original proposed Strassen, we choose to not using them, since the structure of the product is much more complicated, or directly unknown.
@@ 1.2.2 Approximation Theory for Neural Networks
Approximation theory is critical to justify the success of Neural Networks. It studies its density and convergence to the desired map with the objective of answering when a Neural Network is suited for a given job.
@@
The first results in approximation theory for Neural Networks are density theorems over $C^{0}$ and $L^{p}$ spaces. These results are known as universal approximation theorems. The seminal work is [Cyb89], which proved the density of one layer Neural Networks with sigmoidal activation functions in the space of continuous functions equipped with the supremum norm. Generalizations followed, like [HSW89] where a more general set of maps containing Neural Networks is proved to be dense in $C^{0}$, and finally, in [LLPS93] the authors characterize for what activation functions the set of Neural Networks is dense in $L^{p}$ with $p\in(1,\infty)$. Density results are not limited to classic Lebesgue spaces. In fact, the author of [Hor91] gives conditions on the activation function ($m$ times continuously derivable and bounded) for the density in Sobolev spaces over compact and finite measures. As for Besov spaces, its relations with the approximation classes of Neural Network is reflected in [GKNV22]. However, in these papers there is no computation on the number of weights required to approximate a given function.
@@
In order to give upper bounds for the width and depth of Neural Network, the current strategy is to approximate smooth functions and get an estimation for those functions only. Examples of this are given in the papers [Bar93, Mha96, Yar17, PV18, Han19, GKP20, OPS20, GR21, Sie23]. One of the first is [Bar93], where the author gives a relation between the number of weights of a one layer Neural Network and its proximity in $L^{2}$ norm to functions with an integrability condition on its Fourier Transform. In [GKP20] and [GR21], the authors give upper bounds on the number of weights and layers for approximating in the Sobolev space $W^{s,p}$ by Neural Networks a function in $W^{n,p}$ with $n>s$. For [Mha96], [Yar17], and [Sie23], the $L^{p}$ proximity of functions in $W^{s,p}$ with Neural Networks is expressed in terms of number of weights and layers, and the Sobolev norm. In [OPS20], the authors do the same but for the Besov spaces. For $L^{p}$ norms, the authors of [PV18] shows the proximity of continuously differentiable to Neural Networks while tracking the necessary number of parameters. Finally, given a continuous function, in [EHJ17] it is shown that there exists a sequence of Neural Networks with known number of parameters converging in the supreme norm to that function. Check the survey [GRK22] for a general overview of this topic.
@@
From a more computational perspective and non density results, the paper [DLL+19] proves that the most widely used algorithm for training Neural Networks, gradient descent, achieves the desired result as long as the activation function is smooth and Lipschitz. On the same topic, [AZLS19] guarantees, for the same algorithm and for neural networks sufficiently large, convergence as well as there are estimates on its time cost.
@@ 1.2.3 Neural Networks and Multiplication
For scalar multiplication, in [Yar17] and [PV18] a $\operatorname{ReLU}$ Neural Network capable of approximating the product of any pair of scalars in an interval is described. Then, using that Neural Network as building block, the authors of [KPRS22] construct Neural Networks with $\operatorname{ReLU}$ as activation function which approximates the matrix multiplication and inverse operators. This is, as far as we know, the first $\operatorname{ReLU}$ Neural Network to approximate the matrix multiplication or its inverse with upper bounds on the number of weights and layers. For other activation functions like $\varrho(x)=x^{2}$ and $\operatorname{ReLU}^{2}$, those results are obvious while others, even if it is known to exists, there are no upper bounds on the number of parameters for a Neural Network that approximates scalar product operator.
On the reciprocal problem of this work, Strassen-type matrix multiplication algorithms have being obtained using Neural Networks. Those are [Els16, FBH+22] among others. However, there are prior algorithms more optimal. One can find a general review in [MKA24].
Optimization of evaluations and training of Neural Networks or other related architectures of Machine Learning is a huge and ever evolving research topic. Among many other algorithms, the Strassen-type matrix multiplication algorithms are used in this context, we refer for example [ZWWL18, AYK+20, RR18].
@@ 1.2.4 Neural Networks and PDEs
Machine Learning techniques for numerically solving PDEs are a growing research field. The most relevant of this trend is the Physics Informed Neural Networks introduced in [SS18] and [RPK19]. A Physics Informed Neural Network is a Neural Network that approximates the solution of a PDE by minimizing its value to satisfy the equations on a sample of points in the domain. Other example of using Neural Networks as numerical solution of specific PDEs that are not Physics Informed Neural Networks are in [EHJ17, HJKN20].
More recently, the aim of the architectures have shifted from approximating particular solutions of a PDE to approximating its solver operator. One of the first of this kind to appeared is the Deep Operator Net in [LJK19]. Neural Operators are a promising new approach for this task as explained in [KLL+23].
There is also a clash between the classical algorithm for numerically solving PDEs with this new wave of Machine Learning based algorithm. This results on mixing the two approaches in order to mitigate the computational costs of both techniques. Such classic algorithms are the Galerkin method and the Finite Element Method. Combining those methods, which requires an appropriate selection of functions, with Neural Networks we get results such as explained in [HLXZ18, LC20, DL21, ZCL+21, GPR+21]. The idea is to replace in a way the hard to find functions that forms a basis to an trainable family of Neural Networks. Other ideas are exploited in [KPRS22] aiming the mathematical compression instead of computational efficiency. This is not limited either to Neural Networks, for example, the authors of [GZW22] propose Graph Convolutional Networks for this task.
@@ 1.3 Outline
In Section 2, we fix notation, describe the Strassen Algorithm, and define Matrix Neural Networks and related concepts. In Section 3, we construct Matrix Neural Networks that approximates the square and rectangle matrix multiplication operator. In Section 4, we construct a Matrix Neural Network that approximates the inverse of a set of matrices. In Section 5, we explain the use of the previous Matrix Neural Network for numerical solving PDEs. Finally, in Section 6, we compare our results with the ones in [KPRS22].
@@ 2.1 Linear Algebra
In this paper we denote the spaces of matrices of $n_{1}$ rows and $n_{2}$ columns by $\mathbb{R}^{n_{1}\times n_{2}}$. Given $A\in\mathbb{R}^{n_{1}\times n_{2}}$, $i\in\left\{1,\dots,n_{1}\right\}$ and $j\in\left\{1,\dots,n_{2}\right\}$, $A_{i,j}$ denotes the entry of such matrix in its $i$-th row and $j$-th column. Moreover, given $c_{1},c_{2}\in\left\{1,\dots,n_{1}\right\}$ such that $c_{1}\leq c_{2}$ and $c_{3},c_{4}\in\left\{1,\dots,n_{2}\right\}$ such that $c_{3}\leq c_{4}$ we denote by $A_{c_{1}:c_{2},c_{3}:c_{4}}$ its submatrix formed by the rows from $c_{1}$ to $c_{2}$ (including $c_{1}$ and $c_{2}$), and the columns from $c_{3}$ to $c_{4}$ (including $c_{3}$ and $c_{4}$). We also denote by $\operatorname{I}_{n}\in\mathbb{R}^{n\times n}$ the identity matrix with entries $(\operatorname{I}_{n})_{i,j}=\delta_{i,j}$ for all $i,j\in\left\{1,\dots,n\right\}$.
While working with Neural Networks, it is important the notion of matrix concatenation which we define as follows:
@@
In addition, we consider the following notation that is specially useful for describing Strassen algorithm:
@@
Note that we have
@@
Throughout this paper, we consider linear functionals between spaces of matrices. In particular, we consider tensor spaces located in $\mathbb{R}^{(n_{1}\times n_{2})\times(n_{3}\times n_{4})}$. We recall that, given $\mathcal{L}\in\mathbb{R}^{(n_{1}\times n_{2})\times(n_{3}\times n_{4})}$ and $A\in\mathbb{R}^{n_{1}\times n_{2}}$, the product $\mathcal{L}A$ is the matrix in $\mathbb{R}^{n_{3}\times n_{4}}$ given by:
$(\mathcal{L}A)_{k,l}=\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\mathcal{L}_{i,j,k,l}A_{i,j}.$
@@
We will also need to define the following norms:
@@
Let $p\in[1,\infty]$, $A\in\mathbb{R}^{n_{1}\times n_{2}}$ be a matrix and $x\in\mathbb{R}^{n_{1}}$. Then, we denote the $p$-norm of A and $x$ by:
$\left\lvert A\right\rvert_{\infty}\coloneq\begin{cases}\left(\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}\left\lvert A_{i,j}\right\rvert^{p}\right)^{\frac{1}{p}}&\text{ if }p<\infty\\ \sup\limits_{\begin{subarray}{c}i=1,\dots,n_{1}\\ j=1,\dots,n_{2}\end{subarray}}\left\lvert A_{i,j}\right\rvert&\text{ if }p=\infty\end{cases},$ $\left\lvert x\right\rvert_{p}=\begin{cases}\left(\sum\limits_{k=1}^{n_{1}}\left\lvert x_{i}\right\rvert^{p}\right)^{\frac{1}{p}}&\text{ if }p<\infty\\ \sup\limits_{i=1,\dots,n_{1}}\left\lvert x_{i}\right\rvert&\text{ if }p=\infty\end{cases}$
It should not be confused with induced norms even if norms are equivalent in finite dimensional spaces:

### v2

@@ Abstract:
We construct a Neural Network that approximates the matrix multiplication operator for any activation function such that there exists a Neural Network which can approximate the scalar multiplication function. In particular, we use the Strassen algorithm to reduce the number of weights and layers needed for such Neural Networks. This allows us to define another Neural Network for approximating the inverse matrix operator. Also, by relying on the Galerkin method, we apply those Neural Networks to solve parametric elliptic PDEs for a whole set of parameters. Finally, we discuss improvements with respect to the prior results.
@@ 1 Introduction
The Machine Learning explosion has left few research branches untouched, including Numerical Analysis for PDEs. An example are Physics Informed Neural Networks, which provide numerical solutions of PDEs (see [RPK19]). However, this is not the only example: [KPRS22] also constitutes an important breakthrough. There, the authors construct two $\operatorname{ReLU}$ Neural Networks capable of approximating the matrix multiplication and matrix inversion operators while giving an upper bound on the number of weights and layers (see Section 6 for the definition of $\operatorname{ReLU}$ activation function). These Neural Networks are then used in the Galerkin method for solving elliptic PDEs.
The cost of using these methods over conventional ones is hard to justify for solving a single particular PDE. However, they are useful to solve numerically parametric PDEs for a set of parameters. In fact, the use of Machine Learning for parametric PDEs tries to avoid the problems of solving them for each parameter. These problems are essentially the study of too many parameters or a high computing complexity for even a single PDE which both leads to an unreasonable time cost. When solving for a set of parameters at the same time, those problems are mitigated.
In this paper, we focus on parametric elliptic PDEs, which appear in many contexts such as physics and engineering. The problem of describing a physical phenomenon sometimes translates into solving an elliptic PDE. Many examples of this exist, such as the stationary Navier-Stokes, Poisson, Laplace, Helmholtz, and anisotropic heat diffusion equations. These equations explain fluid mechanics, electromagnetism, gravitational fields, and steady-state processes of heat/mass transfers respectively, as shown, for example, in [Pol01, PZ03]. Given their frequency and importance, it is necessary to understand the possible numerical algorithms to efficiently solve these equations.
@@ 1.1 Our Contribution
Motivated by the paper [KPRS22], we seek to understand how Neural Networks can approximate the solutions of parametric PDEs while improving the bounds on the number of parameters via mathematical arguments. For that, we also rely on the Galerkin method combined with the inversion of matrices obtained as the sum of a power series, though in a more precise way. Our contributions are the following ones:
@@
  • Activation functions. In Neural Networks, the activation function plays a major role, as it determines its expressivity. In this paper, we focus on activation functions for which there are Neural Networks that approximate the product operation between two real numbers, as mathematically stated in Hypothesis 1. Clearly, if we cannot approximate such operation, there is no hope of representing the product of matrices, which is needed to apply afterwards the Galerkin method. This set of maps contains $\operatorname{ReLU}$, the activation function considered in [KPRS22].
@@
  • Bounds. The authors of [KPRS22] bounded the number of parameters needed for Neural Networks used for parametric PDEs. To achieve more precise bounds, we use two key ideas. First, we use Strassen algorithm for matrix multiplication, which reduces the number of scalar multiplications needed for matrix multiplication. This allows us to economize the use of the potentially bulky Neural Network that approximates the product of scalars. Secondly, we use a finite sum of powers of matrices to approximate the inverses of matrices appearing in the Galerkin method. Rewriting that sum, we reduce the number of matrix multiplications in comparison to [KPRS22].
@@
    In particular, under the hypothesis on the activation function, we prove two main theorems, which are compared to those in [KPRS22] in Section 6.
@@
    • Theorem 3.4 provides a Neural Network that approximates the matrix multiplication operator. We achieve a smaller size of Neural Network than the one in [KPRS22] with a dependency on the number of rows of the input matrices that is no longer cubic, but of the order of $\log_{2}7$.
    • Theorem 4.6 provides a Neural Network that approximates the matrix inverse operator. Again, we achieve a smaller size of Neural Network than the one in [KPRS22] with a dependency on the number of rows of the input matrices that is no longer cubic, but of the order of $\log_{2}7$.
@@
    As in [KPRS22], in Section 5, we see how Theorem 4.6 is used to numerically solve PDEs with the Galerkin method. The most common application of the Galerkin method, the finite element method, suffers from the curse of dimensionality, which we diminish with the exponent $\log_{2}7$.
@@
  • Notation. Finally, we simplify notation by allowing matrices as input of Neural Networks. This avoids flattening matrices, which overloads the notation.
1.2 State of the art
@@
In this subsection, we discuss the results of the literature related to this paper. The most related work is [KPRS22] since our work is a generalization and improvement of its results, as we have explained before.
@@
For the rest of the literature review, in Section 1.2.1 we explore the matrix multiplication algorithms which optimize the amount of scalar multiplications. We follow in Section 1.2.2 by presenting the approximation theory for Neural Networks, which justifies that the product of scalars can be approximated by Neural Networks with a great variety of activation functions. Then, in Section 1.2.3 we explore the different interactions involving matrix multiplication, Neural Networks and the Strassen algorithm. Finally, in Section 1.2.4 we present the uses that Neural Networks receive to numerically solve PDEs.
@@ 1.2.1 Strassen-Like Algorithms
In [Str69], it is shown that for matrices of size $n\times n$, only $\mathcal{O}(n^{\log_{2}7})$ scalar multiplications are required, instead of $\theta(n^{3})$, which is what the definition suggests.
It is an open problem to find the infimum of $\omega$ such that for every $\varepsilon>0$ there is a $(n\times n)$-matrix multiplication algorithm with $\mathcal{O}(n^{\omega+\varepsilon})$ arithmetic operations. For that purpose a number of more optimal algorithms than Strassen appeared. We would like to highlight two of them: [CW90], which describes an algorithm that was not outperformed for 20 years, and [ADW+25], which is the latest and lowest upper bound. Both use the laser method introduced by Strassen in [Str87] that relates trilinear polynomials to the matrix multiplication tensor. This method, however, has a limitation: [AFLG15] shows that using that method and a variety of its generalizations cannot obtain the bound for $\omega=2$, which is a reasonable target.
Even if those papers present better algorithms than the one in [Str69], as the structure of the product is much more complicated or unknown, we leave it for future work.
@@ 1.2.2 Approximation Theory for Neural Networks
Approximation theory is critical to justify the success of Neural Networks. It studies its density and convergence to the desired function with the objective of answering when a Neural Network is suited for a given job.
@@
The first results in approximation theory for Neural Networks are density theorems over $C^{0}$ and $L^{p}$ spaces, results known as Universal Approximation Theorems. The seminal work is [Cyb89], which proved the density of one layer Neural Networks with sigmoidal activation functions in the space of continuous functions equipped with the supremum norm. Generalizations followed, such as [HSW89], where a more general set of maps containing Neural Networks is proved to be dense in $C^{0}$, and such as [LLPS93], where the authors characterize for what activation functions the set of Neural Networks is dense in $L^{p}$ with $p\in(1,\infty)$. Density results are not limited to classic Lebesgue spaces. In fact, the author of [Hor91] gives conditions on the activation function ($m$ times continuously derivable and bounded) for the density in Sobolev spaces over compact and finite measures. As for Besov spaces, its relations with the approximation classes of Neural Network is reflected in [GKNV22]. However, in these papers, there is no computation on the number of weights required to approximate a given function.
@@
In order to give upper bounds for the width and depth of Neural Network, some smoothness of the function is usually required. Examples of this are given in the papers [Bar93, Mha96, Yar17, PV18, Han19, GKP20, OPS20, GR21, Sie23]. The pioneer is [Bar93], which gives a relation between the number of weights of a one layer Neural Network and its proximity in $L^{2}$ norm to functions with an integrability condition on its Fourier Transform. For [Mha96], [Yar17], and [Sie23], the $L^{p}$ proximity of functions in $W^{s,p}$ with Neural Networks is expressed in terms of the number of weights and layers, and the Sobolev norm. In [OPS20], the authors do the same but for the Besov spaces. For $L^{p}$ norms, the authors of [PV18] shows the proximity of continuously differentiable to Neural Networks while tracking the necessary number of parameters. Moreover, given a continuous function, in [EHJ17] it is shown that there exists a sequence of Neural Networks with known number of parameters converging in the supreme norm to that function. In addition, in [GKP20] and [GR21], the authors provide upper bounds on the number of weights and layers to approximate in the Sobolev space $W^{s,p}$ by Neural Networks a target function which belongs to $W^{n,p}$, for $n>s$. Finally, the survey [GRK22] provides a general overview of this topic.
@@
From a more computational perspective, the paper [DLL+19] proves that the most widely used algorithm for training Neural Networks, gradient descent, achieves the desired result as long as the activation function is smooth and Lipschitz. On the same topic, [AZLS19] guarantees, for the same algorithm and for sufficiently large Neural Networks, convergence, as well as providing estimates on its time cost.
@@ 1.2.3 Neural Networks and Multiplication
For scalar multiplication, in [Yar17] and [PV18] a $\operatorname{ReLU}$ Neural Network capable of approximating the product of any pair of scalars in an interval is described. Then, using that Neural Network as building block, the authors of [KPRS22] construct Neural Networks with $\operatorname{ReLU}$ as activation function which approximates the matrix multiplication and inverse operators. This is, as far as we know, the first explicitly constructed $\operatorname{ReLU}$ Neural Network to approximate matrix multiplication or its inverse with upper bounds on the number of weights and layers. For other activation functions like $\varrho(x)=x^{2}$ and $\operatorname{ReLU}^{2}$, those results are obvious while others, even if it is known to exist, there are no upper bounds on the number of parameters for a Neural Network that approximates scalar multiplication operator.
On the reciprocal problem of this work, Strassen-type matrix multiplication algorithms have been obtained using Neural Networks, by [Els16, FBH+22] among others. Additional examples of those algorithms can be found in the review
[MKA24]. However, it should be noted that the algorithms mentioned in Section 1.2.1 are more efficient asymptotically.
Optimization of evaluations and training of Neural Networks and other related architectures of Machine Learning is a huge and ever-evolving research topic. Among many other algorithms, the Strassen-type matrix multiplication algorithms are used in this context. This has been done from a purely computational perspective, for example, in [RR18, ZWWL18, AYK+20]. We are unaware of theoretical analysis of this kind of algorithms involving Strassen-type matrix multiplications.
@@ 1.2.4 Neural Networks and PDEs
Machine Learning techniques for numerically solving PDEs are a growing research field. The most relevant of this trend is the Physics Informed Neural Networks introduced in [SS18] and [RPK19]. A Physics Informed Neural Network is a Neural Network that approximates the solution of a PDE by minimizing its value to satisfy the equations on a sample of points in the domain. Other examples of using Neural Networks that are not Physics Informed Neural Networks to approximate numerically solutions of specific PDEs are in [EHJ17, HJKN20].
More recently, the aim of architectures has shifted from approximating particular solutions of a PDE to approximating its solver operator. One of the first of this kind to appear is the Deep Operator Net in [LJK19]. In fact, Neural Operators are a promising new approach for this task as explained in [KLL+23].
There is also a clash between the classical algorithm for numerically solving PDEs with this new wave of Machine Learning based algorithm. This results in mixing the two approaches to mitigate the computational costs of both techniques. Such classic algorithms are the Galerkin method and the Finite Element Method. Combining those methods, which requires an appropriate selection of functions, with Neural Networks, we get results such as those in [HLXZ20, LC20, DL21, ZCL+21, GPR+21]. The idea is to replace the hard-to-find functions that form a basis by a trainable family of Neural Networks. Other ideas are exploited in [KPRS22] with the aim of mathematical understanding instead of computational efficiency. Finally, mixing the two approaches is not limited either to Neural Networks. For example, the authors of [GZW22] propose Graph Convolutional Networks for this task.
@@ 1.3 Outline
The outline of the paper is the following: in Section 2, we fix the notation, describe the Strassen Algorithm, and define Matrix Neural Networks and related concepts. In Section 3, we construct Matrix Neural Networks that approximate the matrix multiplication operator. In Section 4, we construct a Matrix Neural Network that approximates the inverse of a set of matrices. In Section 5, we explain the use of the previous Matrix Neural Network for numerically solving PDEs. Finally, in Section 6, we compare our results with those in [KPRS22].
@@ 2.1 Linear Algebra
In this paper, we denote the spaces of matrices of $n_{1}$ rows and $n_{2}$ columns by $\mathbb{R}^{n_{1}\times n_{2}}$. Given $A\in\mathbb{R}^{n_{1}\times n_{2}}$, $i\in\left\{1,\dots,n_{1}\right\}$ and $j\in\left\{1,\dots,n_{2}\right\}$, $A_{i,j}$ denotes the entry in its $i$-th row and $j$-th column. Moreover, given $c_{1},c_{2}\in\left\{1,\dots,n_{1}\right\}$ such that $c_{1}\leq c_{2}$ and $c_{3},c_{4}\in\left\{1,\dots,n_{2}\right\}$ such that $c_{3}\leq c_{4}$ we denote by $A_{c_{1}:c_{2},c_{3}:c_{4}}$ its submatrix formed by the rows from $c_{1}$ to $c_{2}$ (including $c_{1}$ and $c_{2}$), and the columns from $c_{3}$ to $c_{4}$ (including $c_{3}$ and $c_{4}$). We also denote by $\operatorname{I}_{n}\in\mathbb{R}^{n\times n}$ the identity matrix whose entries $(\operatorname{I}_{n})_{i,j}=\delta_{i,j}$ for all $i,j\in\left\{1,\dots,n\right\}$.
While working with Neural Networks, it is important to formalize the notion of matrix concatenation which we define as follows:
@@
In addition, we consider the following notation that is especially useful for describing the Strassen algorithm:
@@
Note that we have:
@@
Throughout this paper, we consider linear functionals between spaces of matrices. In particular, we consider tensors located in $\mathbb{R}^{(n_{1}\times n_{2})\times(n_{3}\times n_{4})}$. We recall that, given $\mathcal{L}\in\mathbb{R}^{(n_{1}\times n_{2})\times(n_{3}\times n_{4})}$ and $A\in\mathbb{R}^{n_{3}\times n_{4}}$, the product $\mathcal{L}A$ is the matrix in $\mathbb{R}^{n_{3}\times n_{4}}$ given by:
$(\mathcal{L}A)_{i,j}=\sum_{l=1}^{n_{4}}\sum_{k=1}^{n_{3}}\mathcal{L}_{i,j,k,l}A_{k,l}.$
In particular, when $n_{1}\times n_{2}=n_{3}\times n_{4}$, we define the identity tensor $\operatorname{id}$, whose entries are:
$(\operatorname{id})_{i,j}=\delta_{i\times j,k\times l}.$
@@
We will also need the following norms:
@@
Let $p\in[1,\infty]$, $A\in\mathbb{R}^{n_{1}\times n_{2}}$ and $x\in\mathbb{R}^{n_{1}}$. Then, we denote the $p$-norm of $A$ and $x$ by:
$\left\lvert A\right\rvert_{p}\coloneq\begin{cases}\left(\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}\left\lvert A_{i,j}\right\rvert^{p}\right)^{\frac{1}{p}}&\text{ if }p<\infty,\\ \sup\limits_{\begin{subarray}{c}i=1,\dots,n_{1}\\ j=1,\dots,n_{2}\end{subarray}}\left\lvert A_{i,j}\right\rvert&\text{ if }p=\infty,\end{cases}$ $\left\lvert x\right\rvert_{p}=\begin{cases}\left(\sum\limits_{k=1}^{n_{1}}\left\lvert x_{i}\right\rvert^{p}\right)^{\frac{1}{p}}&\text{ if }p<\infty,\\ \sup\limits_{i=1,\dots,n_{1}}\left\lvert x_{i}\right\rvert&\text{ if }p=\infty.\end{cases}$
It should not be confused with induced norms, even if norms are equivalent in finite-dimensional spaces:
@@
We recall the following estimates between different norms:
Lemma 2.6.
Let $A\in\mathbb{R}^{n\times n}$. Then,
$\left\lvert A\right\rvert_{\infty}\leq\left\lVert A\right\rVert_{2}\leq n\left\lvert A\right\rvert_{\infty}.$ (2.1)